# Quantum Down Sampling Filter for Variational Auto-encoder

### v1

@@ 6.1 Reconstructed Images:
On the MNIST dataset, the Q-VAE achieved the lowest MSE of 0.0088, demonstrating its ability to produce high-fidelity reconstructions compared to the other models. This is consistent with the Q-VAE’s superior performance across other metrics such as training and testing loss. The quantum-based encoding method enables the Q-VAE to extract richer feature representations, leading to more accurate reconstructions. The quantum model’s ability to capture complex dependencies between pixels allows it to outperform classical approaches, even when dealing with highly structured images like those in the MNIST dataset. The C-VAE, with an MSE of 0.0093, showed slightly higher error compared to the Q-VAE, indicating that while it performs well, its feature extraction capabilities are not as advanced as those of the quantum model. Despite its relatively low MSE, the classical model’s performance is constrained by the limitations of traditional neural network-based encoding and decoding methods, which cannot capture the full complexity of the data as efficiently as quantum circuits. The CDP-VAE, with an MSE of 0.0092, performs little better then C-VAE but still didn’t perform better then Q-VAE. USPS images reconstructed from 3 different VAE models have been shown in Figure 11.
@@
While this score indicates that the model can generate images with reasonable fidelity (i.e., the generated images resemble real USPS digits), it lags behind the performance of the Q-VAE. The quantum model’s ability to leverage more complex transformations in the latent space for image reconstruction enables it to capture more intricate patterns and feature representations in the data. This allows the Q-VAE to achieve lower FID scores (indicating higher image fidelity) compared to the CDP-VAE and C-VAE as seen in the results for the USPS dataset. For MSE, the Q-VAE again outperformed the other models, achieving a score of 0.0058. This lower error indicates that the quantum model is particularly effective at capturing the more complex, high-resolution patterns in the USPS images. The Q-VAE’s quantum-enhanced encoding allows it to extract intricate features from the grayscale images, leading to more accurate reconstructions. Its performance on USPS highlights the potential of quantum circuits in improving image generation tasks, especially for datasets with more complex structures compared to MNIST. The C-VAE, with an MSE of 0.0070, performed worse than the Q-VAE and CDP-VAE on USPS, reflecting the same trend observed in MNIST. While still a competitive model, the C-VAE’s ability to reconstruct images is limited by its classical encoding mechanism. The model performs well, but its lack of advanced feature extraction techniques prevents it from achieving the same reconstruction quality as the Q-VAE. The CDP-VAE with an MSE of 0.0065, performance is better than the C-VAE on MNIST, it still lags behind both the Q-VAE on USPS. While reconstructed images we can see Q-VAE is reconstructing very well competing with C-VAE and CDP-VAE.
@@ 6.2 New Generated Images:
The performance of the three models in handling random noise varies considerably. The generated images presented were captured at epoch 200. The Classical VAE requires more time to converge, necessitating longer training. Therefore, we evaluate the model at epoch 200 to assess its performance after sufficient training. While initial epochs Q-VAE has shown fast convergence at epoch 20, while other models were unable to generate digits even at epoch 20 .The Q-VAE outperforms the other models, achieving an FID score of 78.7 at epoch 200. This suggests that the Q-VAE is more resilient to noise, as it is capable of generating higher-quality images even when starting from random noise inputs. In contrast, the C-VAE struggles with random noise, resulting in a much higher FID score of 94.4, indicating poor performance in reconstructing images from noisy inputs. The CDP-VAE also faces challenges with noise but performs slightly better than the C-VAE, with an FID score of 93.3. While both the C-VAE and CDP-VAE show limitations in noise handling, the Q-VAE demonstrates a notable advantage in generating more realistic images in the presence of noise. MNIST generated images from noise, have been shown below in Figure 12:
@@
Interestingly, Q-VAE have demonstrated better performance in similar tasks, as reflected in improved FID scores. Q-VAE, in particular, have shown a remarkable ability to generate high-quality images from noise These operations are highly effective at capturing complex patterns and correlations in high-resolution data, enabling more detailed and coherent outputs. Figure 13 shows how well Quantum can generate new images from noise while using only quantum superposition. Quantum advantage enables the model to effectively capture complex patterns from random noise, producing high-quality, realistic images even in challenging condition [16] This highlights the advantage of quantum-based encoding in generating high-quality images, particularly in tasks that require nuanced and high-resolution data representations, such as image reconstruction or generation.
6.2.1 Analysis of Results: Quantum VAE Performance
The quantum encoding in the Q-VAE allows it to produce better results than classical models, particularly in noisy conditions. The use of Quantum states representation helps the Q-VAE achieve more robust image reconstructions. This advantage is especially evident in the FID scores under random noise conditions, where the quantum model maintains high fidelity even when the input is random noise. The MSE analysis confirms that the Q-VAE consistently outperforms both the C-VAE and CDP-VAE on both the MNIST and USPS datasets. The quantum-enhanced feature extraction capability of the Q-VAE leads to the most accurate image reconstructions, as evidenced by the lowest MSE values across both datasets. While the C-VAE performs well, its classical encoding mechanisms are not as effective in capturing complex data distributions as the quantum model. The CDP-VAE, due to its simpler approach, still outperforms C-VAE. Furthermore, the quantum circuit’s probabilistic nature may act as a form of regularization, leading to better generalization, as demonstrated by the lower FID scores in both latent space and noise-based evaluations.
@@ 6.2.2 Trainable Parameters
The results indicates that our proposed model Q-VAE and its similar approach in classical, CDP-VAE, both have shown significant improvement over C-VAE architectures in terms of parameter efficiency. Specifically, the C-VAE has a total of 407,377 trainable parameters, with 299,424 parameters dedicated to the encoder and 107,953 parameters assigned to the decoder. In contrast, the Quantum VAE significantly reduces the number of trainable parameters, totaling only 144,977 parameters. This is achieved by the quantum model’s more compact encoder, which contains only 37,024 parameters, a dramatic reduction compared to the classical VAE. This reduction highlights the efficiency of the quantum approach in learning and representing complex features with a smaller number of parameters, enabling better latent space compression. Despite this reduction in parameters, the decoder structure remains similar between the quantum and classical models, both utilizing 107,953 parameters. This consistency in the decoder ensures that the Quantum VAE maintains comparable image reconstruction capabilities to the classical model, while benefiting from the quantum encoder’s enhanced efficiency. The lower parameter count in the QVAE encoder illustrates the potential of quantum circuits to offer a more resource-efficient alternative for training and generating high-quality image reconstructions. These findings underscore the practical benefits of quantum computing in optimizing VAE architectures, especially in tasks that require high-fidelity image generation and accurate reconstruction. While classical VAEs remain a strong choice for many machine learning applications, the Q-VAE shows promising advancements, particularly in fields where image quality and computational efficiency are paramount. Potential applications of quantum VAEs include digital imaging, healthcare diagnostics, and creative arts, where high-quality image generation is critical. However, despite these advantages, the adoption and effectiveness of Q-VAE still face challenges, including issues related to the scalability of quantum hardware and the complexity of quantum algorithms. Continued research and development in quantum machine learning are essential for unlocking the full potential of quantum-enhanced models, pushing the boundaries of what’s achievable in generative modeling and real-world applications.
@@ 7 Conclusion
In this study, we introduce the Quantum Down-sampling Filter for Variational Auto-encoder (Q-VAE), a novel model that integrates quantum computing into the encoder component of a traditional VAE. Our experimental results demonstrate that the Q-VAE offers significant advantages over classical VAE architectures across various performance metrics. Notably, the Q-VAE consistently outperforms the classical VAE, achieving lower FID scores, which indicate superior image fidelity and the ability to generate higher-quality reconstructions. The quantum-enhanced encoder captures more intricate features of images, leading to improved generative performance. We evaluated the model on digit datasets such as MNIST and USPS, initially up-scaled to 32x32 for consistency. For a fair comparison, we also tested our quantum-enhanced model alongside a classical counterpart and a model mimicking classical behavior, referred to as Classical Direct Passing (CDP-VAE). One more key finding of this study is the significant reduction in the number of trainable parameters in the Q-VAE’s and CDP-VAE’s encoder. With only 37,024 parameters, the Q-VAE dramatically lowers computational complexity compared to the classical VAE, which contains over 407,000 trainable parameters. This reduction highlights the efficiency of quantum-based encoding methods, offering the potential for more scalable and resource-efficient models. The results highlight the practical advantages of incorporating quantum computing into VAEs, particularly for tasks requiring high-quality image generation and feature extraction. The Q-VAE not only improves image quality but also enhances computational efficiency, making it a promising candidate for future applications in areas such as computer vision, data synthesis, and beyond. These findings suggest that further exploration of quantum-enhanced machine learning techniques could unlock new possibilities in generative modeling and image reconstruction, offering both performance improvements and computational savings.

### v2

@@ 6.1 Reconstructed Images:
USPS images reconstructed from 3 different VAE models have been shown in Figure 10.
@@
While this score indicates that the model can generate images with reasonable fidelity (i.e., the generated images resemble real USPS digits), it lags behind the performance of the Q-VAE. The quantum model’s ability to leverage more complex transformations in the latent space for image reconstruction enables it to capture more intricate patterns and feature representations in the data. This allows the Q-VAE to achieve lower FID scores (indicating higher image fidelity) compared to the CDP-VAE and C-VAE as seen in the results for the USPS dataset. For MSE, the Q-VAE again outperformed the other models, achieving a score of 0.0058. This lower error indicates that the quantum model is particularly effective at capturing the more complex, medium-resolution patterns in the USPS images. The Q-VAE’s quantum-enhanced encoding allows it to extract intricate features from the grayscale images, leading to more accurate reconstructions. Its performance on USPS highlights the potential of quantum circuits in improving image generation tasks, especially for datasets with more complex structures compared to MNIST. The C-VAE, with an MSE of 0.0070, performed worse than the Q-VAE and CDP-VAE on USPS, reflecting the same trend observed in MNIST. While still a competitive model, the C-VAE’s ability to reconstruct images is limited by its classical encoding mechanism. The model performs well, but its lack of advanced feature extraction techniques prevents it from achieving the same reconstruction quality as the Q-VAE. The CDP-VAE with an MSE of 0.0065, performance is better than the C-VAE on MNIST, it still lags behind both the Q-VAE on USPS. While reconstructed images we can see Q-VAE is reconstructing very well competing with C-VAE and CDP-VAE.
@@ 6.2 New Generated Images:
The performance of the three models in handling random noise varies considerably. The generated images presented were captured at epoch 200. The Classical VAE requires more time to converge, necessitating longer training. Therefore, we evaluate the model at epoch 200 to assess its performance after sufficient training. While initial epochs Q-VAE has shown fast convergence at epoch 20, while other models were unable to generate digits even at epoch 20 .The Q-VAE outperforms the other models, achieving an FID score of 78.7 at epoch 200. This suggests that the Q-VAE is more resilient to noise, as it is capable of generating better-quality images even when starting from random noise inputs. In contrast, the C-VAE struggles with random noise, resulting in a much higher FID score of 94.4, indicating poor performance in reconstructing images from noisy inputs. The CDP-VAE also faces challenges with noise but performs slightly better than the C-VAE, with an FID score of 93.3. While both the C-VAE and CDP-VAE show limitations in noise handling, the Q-VAE demonstrates a notable advantage in generating more realistic images in the presence of noise. MNIST generated images from noise, have been shown below in Figure 11:
@@
Interestingly, Q-VAE have demonstrated better performance in similar tasks, as reflected in improved FID scores. Q-VAE, in particular, have shown a remarkable ability to generate better-quality images from noise These operations are highly effective at capturing complex patterns and correlations in medium-resolution data, enabling more detailed and coherent outputs. Figure 12 shows how well Quantum can generate new images from noise while using only quantum superposition. The quantum advantage enables the model to effectively capture complex patterns from random noise, producing high-quality, leading to more realistic images [17] This highlights the advantage of quantum-based encoding in generating better-quality images, particularly in tasks that require nuanced and medium-resolution data representations, such as image reconstruction or generation.
@@ 6.2.1 Trainable Parameters
The results indicates that our proposed model Q-VAE and its similar approach in classical, CDP-VAE, both have shown significant improvement over C-VAE architectures in terms of parameter efficiency. Specifically, the C-VAE has a total of 407,377 trainable parameters, with 299,424 parameters dedicated to the encoder and 107,953 parameters assigned to the decoder. In contrast, the Quantum VAE significantly reduces the number of trainable parameters, totaling only 144,977 parameters. This is achieved by the quantum model’s more compact encoder, which contains only 37,024 parameters, a dramatic reduction compared to the classical VAE. This reduction highlights the efficiency of the quantum approach in learning and representing complex features with a smaller number of parameters, enabling better latent space compression.
Despite this reduction in parameters, the decoder structure remains similar between the quantum and classical models, both utilizing 107,953 parameters. This consistency in the decoder ensures that the Quantum VAE maintains comparable image reconstruction capabilities to the classical model, while benefiting from the quantum encoder’s enhanced efficiency. The lower parameter count in the QVAE encoder illustrates the potential of quantum circuits to offer a more resource-efficient alternative for training and generating better-quality image reconstructions. These findings underscore the practical benefits of quantum computing in optimizing VAE architectures, especially in tasks that require high-fidelity image generation and accurate reconstruction. While classical VAEs remain a strong choice for many machine learning applications, the Q-VAE shows promising advancements, particularly in fields where image quality and computational efficiency are paramount. Potential applications of quantum VAEs include digital imaging, healthcare diagnostics, and creative arts, where high-quality image generation is critical. However, despite these advantages, the adoption and effectiveness of Q-VAE still face challenges, including issues related to the scalability of quantum hardware and the complexity of quantum algorithms. Continued research and development in quantum machine learning are essential for unlocking the full potential of quantum-enhanced models, pushing the boundaries of what is achievable in generative modeling and real-world applications.
@@ 7 Conclusion
In this study, we introduce the Quantum Down-sampling Filter for Variational Auto-encoder (Q-VAE), a novel model that integrates quantum computing into the encoder component of a traditional VAE. Our experimental results demonstrate that Q-VAE offers significant advantages over classical VAE architectures across various performance metrics. In particular, Q-VAE consistently outperforms classical VAE, achieving lower FID scores, indicating superior image fidelity and the ability to generate better-quality reconstructions. The quantum-enhanced encoder captures more intricate features of the images, leading to improved generative performance. We evaluated the model on digit datasets such as MNIST and USPS, initially up-scaled to 32x32 for consistency. For a fair comparison, we also tested our quantum-enhanced model alongside a classical counterpart and a model mimicking classical behavior, referred to as Classical Direct Passing (CDP-VAE). One more key finding of this study is the significant reduction in the number of trainable parameters in the Q-VAE’s and CDP-VAE’s encoder. With only 37,024 parameters, the Q-VAE dramatically reduces computational complexity compared to the classical VAE, which contains over 407,000 trainable parameters. This reduction highlights the efficiency of quantum-based encoding methods, which offers the potential for more scalable and resource-efficient models. The results highlight the practical advantages of incorporating quantum computing into VAEs, particularly for tasks that require high-quality image generation and feature extraction. The Q-VAE not only improves image quality, but also enhances computational efficiency, making it a promising candidate for future applications in areas such as computer vision, data synthesis, and beyond. These findings suggest that further exploration of quantum-enhanced machine learning techniques could unlock new possibilities in generative modeling and image reconstruction, offering both performance improvements and computational savings.